OpenStudy (anonymous):

Perform the multiplication on the 12-hour clock. 11 x 3 i belive its 9

OpenStudy (anonymous):

yes it's 9

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Find the additive inverse of 5 on the 12-hour clock. i cant figre this out

OpenStudy (anonymous):

i think 10

OpenStudy (anonymous):

im going over y answers

OpenStudy (anonymous):

The additive inverse of 5 is the number that when added to 5 will give you the identity in the mod 12 system. So you are solving 5+x=0 mod 12 you get x=-5 but x=-5+12=7 mod 12 so your answer is 7 because when you add 5+7=12 or 0 mod 12 :-)